# Observer-robust energy condition verification for warp drive spacetimes

### 03 Mar 2026 v2

@@ 4.3 Velocity dependence
Figure 6 shows the spatial distribution of missed violations at each velocity. The central result is that the discrepancy between Eulerian and observer-robust analyses is systematic and persistent across the full velocity range, with the strongest effect in the DEC and WEC conditions rather than the NEC.
@@ 5.1 Tidal forces
Figure 8: Tidal eigenvalue evolution along a radial timelike geodesic through the Alcubierre bubble ($v_{s}=0.5$, $R=1$, $\sigma=8$; ODE integration via Tsit5; see Table 1). The three eigenvalues of the tidal tensor $K^{\mu}{}_{\rho}$ are plotted as a function of proper time. Sharp tidal forces at the bubble wall indicate significant stretching and compression.
@@ 5.2 Blueshift analysis
Figure 9: Photon frequency ratio along a null geodesic through the Alcubierre bubble ($v_{s}=0.5$, $R=1$, $\sigma=8$; ODE integration via Tsit5; see Table 1). The blueshift at the leading edge and redshift at the trailing edge reflect the asymmetric spacetime distortion.
@@ 7 Resolution stability
Table 2 characterizes the wall resolution for each metric at $50^{3}$. Rodal is the only metric that comfortably exceeds the 4-cell resolution threshold ($\sim\!6$ cells); Alcubierre, VdB, and Natário are marginally resolved ($\sim\!1.4$ cells), Lentz is severely under-resolved ($\sim\!0.02$ cells), and WarpShell concentrates its thin-shell transition within $\lesssim\!1$ cell. Autodiff computes exact curvature at each grid point regardless of wall resolution. To verify that the headline missed-violation fractions are not resolution artifacts, we run the same three-resolution sweep ($25^{3}$$50^{3}$$100^{3}$) for the Rodal ($R=100$, $\sigma=0.03$) and WarpShell ($R_{1}=0.5$, $R_{2}=1.0$) metrics. At $100^{3}$ we use Eulerian-only evaluation due to the prohibitive cost of full observer optimization at $10^{6}$ points. The minimum NEC margin converges at second order for Rodal: extrapolation gives $2.58\times 10^{-6}$ ($50^{3}$ value: $2.62\times 10^{-6}$, within 1.5%). For WarpShell, the Eulerian minimum NEC margin increases in absolute magnitude with resolution ($2.5\times 10^{25}$ at $25^{3}$, $2.3\times 10^{33}$ at $50^{3}$, $2.1\times 10^{34}$ at $100^{3}$), reflecting the shell boundary extremes that are better resolved at finer grids; Richardson extrapolation yields a negative convergence order ($p\approx-3.0$), so we report per-resolution values rather than an extrapolated limit. The Rodal integrated violation does not converge monotonically, consistent with the broad Rodal wall ($\sigma=0.03$, 10–90% transition width $73.24$ coordinate units spanning $\sim\!6.0$ grid cells; Table 2), which is marginally resolved at the $50^{3}$ grid but whose small gradients require higher resolution to stabilize extensive (volume-integrated) quantities.
@@
Table 6: Rodal missed-fraction stability across grid resolutions ($v_{s}=0.5$, $R=100$, $\sigma=0.03$, $N_{\text{starts}}=8$). All points are Type I at $N\geq 50$; truth is the algebraic slack. Missed fractions vary by <2<2 percentage points across a $64\times$ increase in grid volume.
@@ 7.2 Optimization stability
To assess observer optimizer reliability, we examine per-point convergence diagnostics for the WarpShell metric at $v_{s}=0.5$ (the only metric for which per-point convergence data was recorded in the current analysis run). The NEC optimizer converges at 99.8% of grid points with a median of 2 iterations; 0.23% of points reach the maximum iteration limit of 256. The WEC and SEC optimizers converge at 99.7% with 0.31% hitting the iteration limit, while the DEC converges at 99.6% with 0.39% at the limit. Across all four conditions, the median iteration count is 2 and the mean is below 3, indicating that the optimization landscape is well-conditioned at the vast majority of grid points. The slightly lower DEC convergence rate is consistent with the DEC objective combining flux causality, future-directedness, and energy density, a more complex landscape than the single-contraction NEC or WEC objectives. These results confirm that the multi-start observer optimization procedure converges reliably, with fewer than 0.4% of points requiring the full iteration budget.
@@ 7.3 Classification breakdown and conditional miss rate
Table 8: Hawking–Ellis type breakdown at $v_{s}=0.5$ ($50^{3}$ grid; Schwarzschild at $v_{s}=0$). Metrics with steep shape functions (Natário, VdB) show slightly higher non-Type I fractions due to numerically split degenerate eigenvalues at transition zones.
Metric % Type I % Type II % Type IV max $|\mathrm{Im}\,\lambda|/s$
Rodal 100.0 0.0 0.0 0
Schwarzschild 100.0 0.0 0.0 $7.3\times 10^{-19}$
Lentz 99.9 0.0 0.1 $2.8\times 10^{-1}$
Alcubierre 98.0 0.0 2.0 $2.7\times 10^{-1}$
Natário 97.5 0.0 2.5 $1.7\times 10^{0}$
VdB 98.4 0.0 1.6 $2.2\times 10^{-1}$
WarpShell 99.6 0.3 0.1 $1.4\times 10^{34}$
@@ 8 Discussion
The central result of this work is that single-frame energy condition evaluation can systematically underestimate the extent of energy condition violations in warp drive spacetimes, though the severity depends strongly on the metric geometry. For the Rodal metric, the Eulerian DEC analysis fails to detect violations at over 28% of all grid points, and the WEC figure exceeds 15%. For WarpShell (regularized implementation), missed WEC violations are below $0.1\%$ of the grid. For the Alcubierre and Natário metrics, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis at the tested parameters (missed =0=0 at the sampled resolution). Lentz shows zero missed NEC/WEC points (though nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; Table 2); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$, DEC $0.3\%$ at $v_{s}=0.5$; Figures 34), and non-trivial SEC misses persist, most notably $1.2\%$ for Van Den Broeck (conditional miss rate $17.5\%$; Table 3). Even for these metrics the $\zeta_{\max}$-capped violation severity can be much larger (Section 2.4, Interpretation caveat): for the Alcubierre WEC at $\zeta_{\max}=5$, the capped minimum is $\sim\!-3450$ compared to 0.038-0.038 for the Eulerian frame. These severity ratios are $\zeta_{\max}$-dependent and grow without bound; single-frame analysis can underestimate the magnitude of violations even when it correctly identifies their spatial extent.
@@
For the Alcubierre metric at $v_{s}=0.5$, $R=1$, $\sigma=8$ (Figure 5), we observe empirically that the worst-case WEC observers are predominantly Lorentz-boosted along the direction of bubble propagation. This is physically plausible: boosted observers see a different effective stress-energy distribution, and the Eulerian frame, adapted to the spatial hypersurface, is not aligned with the worst-case direction. For the Rodal DEC, the alignment angle between the optimizer’s boost direction and the DEC-worst eigenvector spatial direction has a median of $82^{\circ}$ across all DEC-violating Type I points at four velocities (Figure 12), with only 0.90.9% of points having alignment below $10^{\circ}$. This confirms that the worst-case DEC observers are overwhelmingly non-Eulerian and that single-frame analysis misses violations precisely because it fails to probe these nearly orthogonal directions. The effect is condition-dependent: the WEC and DEC are more sensitive to observer choice than the NEC, because they involve timelike observers (parameterized by rapidity) rather than null directions alone.
@@
Any claim that a warp metric satisfies an energy condition must specify which observers were tested and with what coverage. The original claim that “the Lentz metric has positive energy” was shown by Celmaster and Rubin [21] to be incorrect even for Eulerian observers. More generally, our results for Rodal and WarpShell demonstrate that single-frame analysis can miss substantial violation fractions, while for other metrics the Eulerian frame detects all violation points. Continuous observer optimization, as implemented in warpax, quantifies both the spatial extent of missed violations and the severity increase for non-Eulerian observers.
@@
WarpFactory [4, 5, 24] samples $\sim\!1000$ observer directions and velocities at each grid point and takes the minimum energy condition value over the sample set. This discrete-sampling approach correctly identifies observer dependence, but its accuracy depends on the sampling density: narrow violation cones in the observer manifold may be missed if no sample falls within them. A Fibonacci-lattice benchmark (Figure 13) confirms this quantitatively: even $10^{4}$ deterministic samples per point saturate at $\sim\!93$% DEC detection for the Rodal metric, with a sampled minimum margin three orders of magnitude looser than the observer-optimized result. warpax complements sampling with continuous observer optimization (Section 2.4), which targets the worst-case observer via gradient descent with multi-start initialization. The two approaches are complementary: WarpFactory provides a well-tested MATLAB implementation for rapid exploration, while warpax provides gradient-based observer optimization, autodiff curvature, and Hawking–Ellis classification in a GPU-accelerated Python ecosystem. Our Eulerian-baseline NEC/WEC results match WarpFactory’s published values for the Alcubierre metric at matched parameters; residual $\sim\!1$% relative differences are consistent with autodiff vs. finite-difference curvature computation.
@@
The observer optimization is the dominant computational cost. For a single metric at $50^{3}$ resolution with $N_{\text{starts}}=8$ optimizer runs per point, the full analysis takes on the order of minutes on GPU hardware (compared to seconds for the Eulerian-only analysis). For Type I points, the Hawking–Ellis classification provides exact algebraic slacks (for the implemented regularized metric) that bypass the optimization entirely for violation detection; the optimizer cost is then justified primarily by worst-observer identification.
@@
To validate that observer optimization provides genuine value beyond dense observer sampling, we evaluate detection rates at the Eulerian-missed violation points for each metric. For the Rodal NEC and WEC, dense sampling at $10^{4}$ random directions detects 100% of the violations that the optimizer finds. However, for the Rodal DEC, sampling plateaus at only 54% detection even at $10^{4}$ directions, while in our tests the optimizer recovered 100% of the algebraic Type I violation set, a clear case where the violation cone in observer space is too narrow for random sampling to cover reliably. These results demonstrate that while dense sampling suffices for NEC and WEC in most cases, gradient-based optimization substantially outperforms random sampling for DEC analysis, where the violation cone in observer space can be too narrow for random sampling to cover reliably. For Type I points, the algebraic eigenvalue check provides exact DEC truth (for the implemented regularized metric) without any observer search; the optimization advantage is relevant for analyses that rely on observer sampling. Table 10 summarizes these detection rates.
@@
Our analysis is restricted to classical energy conditions. Quantum energy inequalities [25, 26] provide weaker but more physical constraints on the stress-energy, and their systematic evaluation over observers remains an open problem. Additionally, our metrics are parameterized by a finite set of shape parameters ($v_{s}$, $R$, $\sigma$), and the results may not generalize to metrics outside this family.
@@
Several extensions are natural: (i) metric-space optimization, where the shape function $f(r_{s})$ itself is optimized to minimize energy condition violations; (ii) integration of quantum inequality bounds [25]; (iii) application to the broader landscape of warp drive metrics, including the Bobrick–Martire classification [22]; and (iv) extension to non-vacuum backgrounds (e.g., cosmological spacetimes).
@@ 9 Conclusion
The author acknowledges support from the VinUni Center for Environmental Intelligence (CEI). The author thanks J. Rodal for clarifications regarding frame conventions, the $r_{s}\to 0$ limit of the irrotational angular profile, and the Cartesian form of the shift vector. The author declares no competing interests.
@@ Appendix D Metric definitions
$f(r_{s})=\frac{\tanh\bigl[\sigma(r_{s}+R)\bigr]-\tanh\bigl[\sigma(r_{s}-R)\bigr]}{2\,\tanh(\sigma R)}\,$, (27)
with $f(0)=1$ and $f(r_{s})\to 0$ for $r_{s}\gg R$.
@@
$G(r_{s})=1-\frac{2\,r_{s}\sigma\sinh(R\sigma)+\cosh(R\sigma)\bigl[\ln\cosh\sigma(r_{s}{-}R)-\ln\cosh\sigma(r_{s}{+}R)\bigr]}{2\,r_{s}\sigma\sinh(R\sigma)}\,.$ (31)
@@
Although equation (31) contains an apparent $1/r_{s}$ factor, this is a removable $0/0$ form arising from the spherical-basis representation. Defining $\Delta(r_{s})\equiv\ln\cosh\bigl[\sigma(r_{s}{-}R)\bigr]-\ln\cosh\bigl[\sigma(r_{s}{+}R)\bigr]$, one finds $\Delta^{\prime}(0)=-2\sigma\tanh(\sigma R)$, whence $\lim_{r_{s}\to 0}G(r_{s})=1$. The irrotational construction is therefore regular at the bubble center; the $r_{s}\mapsto\sqrt{r_{s}^{2}+\varepsilon^{2}}$ substitution used in our implementation is a numerical stability device for automatic differentiation, not a physical repair.

### 27 Apr 2026 v3

@@ 4.3 Velocity dependence
Figure 6 shows the spatial distribution of missed violations at each velocity. Our central finding is that the discrepancy between Eulerian and observer-robust analyses is systematic and persistent across the full velocity range, with the strongest effect in the DEC and WEC conditions rather than the NEC.
@@ 5.1 Tidal forces
Figure 8: Tidal eigenvalue evolution along a radial timelike geodesic through the Alcubierre bubble ($v_{s}=0.5$, $R_{b}=1$, $\sigma=8$; ODE integration via Tsit5; see Table 1). The three eigenvalues of the tidal tensor $K^{\mu}{}_{\rho}$ are plotted as a function of proper time. Sharp tidal forces at the bubble wall indicate significant stretching and compression.
@@ 5.2 Blueshift analysis
Figure 9: Photon frequency ratio along a null geodesic through the Alcubierre bubble ($v_{s}=0.5$, $R_{b}=1$, $\sigma=8$; ODE integration via Tsit5; see Table 1). The blueshift at the leading edge and redshift at the trailing edge reflect the asymmetric spacetime distortion.
@@ 7 Resolution stability
Table 3 characterizes the wall resolution for each metric at $50^{3}$. Rodal is the only metric that comfortably exceeds the 4-cell resolution threshold ($\sim\!6$ cells); Alcubierre, VdB, and Natário are marginally resolved ($\sim\!1.4$ cells), Lentz is severely under-resolved ($\sim\!0.02$ cells), and WarpShell concentrates its thin-shell transition within $\lesssim\!1$ cell. Autodiff computes exact curvature at each grid point regardless of wall resolution. To verify that the headline missed-violation fractions are not resolution artifacts, we run the same three-resolution sweep ($25^{3}$$50^{3}$$100^{3}$) for the Rodal ($R_{b}=100$, $\sigma=0.03$) and WarpShell ($R_{1}=0.5$, $R_{2}=1.0$) metrics. At $100^{3}$ we use Eulerian-only evaluation due to the prohibitive cost of full observer optimization at $10^{6}$ points. The minimum NEC margin converges at second order for Rodal: extrapolation gives $2.58\times 10^{-6}$ ($50^{3}$ value: $2.62\times 10^{-6}$, within 1.5%). For WarpShell, the Eulerian minimum NEC margin increases in absolute magnitude with resolution ($2.5\times 10^{25}$ at $25^{3}$, $2.3\times 10^{33}$ at $50^{3}$, $2.1\times 10^{34}$ at $100^{3}$), reflecting the shell boundary extremes that are better resolved at finer grids; Richardson extrapolation yields a negative convergence order ($p\approx-3.0$), so we report per-resolution values rather than an extrapolated limit. The Rodal integrated violation does not converge monotonically, consistent with the broad Rodal wall ($\sigma=0.03$, 10–90% transition width $73.24$ coordinate units spanning $\sim\!6.0$ grid cells; Table 3), which is marginally resolved at the $50^{3}$ grid but whose small gradients require higher resolution to stabilize extensive (volume-integrated) quantities.
@@
Table 7: Rodal missed-fraction stability across grid resolutions ($v_{s}=0.5$, $R_{b}=100$, $\sigma=0.03$, $N_{\text{starts}}=8$). All points are Type I at $N\geq 50$; truth is the algebraic slack. Missed fractions vary by <2<2 percentage points across a $64\times$ increase in grid volume.
@@ 7.2 Optimization stability
To assess observer optimizer reliability, we examine per-point convergence diagnostics for the WarpShell metric at $v_{s}=0.5$ (the only metric for which per-point convergence data was recorded in the current analysis run). The NEC optimizer converges at 99.8% of grid points with a median of 2 iterations; 0.23% of points reach the maximum iteration limit of 256. The WEC and SEC optimizers converge at 99.7% with 0.31% hitting the iteration limit, while the DEC converges at 99.6% with 0.39% at the limit. Across all four conditions, the median iteration count is 2 and the mean is below 3, indicating that the optimization problem is well-conditioned at the vast majority of grid points. The slightly lower DEC convergence rate is consistent with the DEC objective combining flux causality, future-directedness, and the observer-contracted energy density $T_{ab}\,u^{a}u^{b}$, a more complex objective than the single-contraction NEC or WEC objectives. These results confirm that the multi-start observer optimization procedure converges reliably, with fewer than 0.4% of points requiring the full iteration budget.
@@ 7.3 Classification breakdown and conditional miss rate
Table 9: Hawking–Ellis type breakdown at $v_{s}=0.5$ ($50^{3}$ grid; Schwarzschild at $v_{s}=0$). Full-grid columns cover all 125,000125{,}000 points; wall-restricted columns filter to the active warp-wall region $f\in[0.1,0.9]$. Metrics with steep shape functions (Natário, VdB) show slightly higher non-Type I fractions due to numerically split degenerate eigenvalues at transition zones.
Metric % Type I % Type II % Type IV Wall % Type I Wall % Type IV max $|\mathrm{Im}\,\lambda|/s$
Alcubierre 98.0 0.0 2.0 1.9 98.1 $2.7\times 10^{-1}$
Natário 97.5 0.0 2.5 9.6 90.4 $1.7\times 10^{0}$
Van Den Broeck 98.4 0.0 1.6 15.4 84.6 $2.2\times 10^{-1}$
Rodal 100.0 0.0 0.0 100.0 0.0 0
Lentz 99.9 0.0 0.1 0.0 100.0 $2.8\times 10^{-1}$
WarpShell 99.8 0.1 0.1 53.6 12.5 $1.4\times 10^{34}$
Schwarzschild 100.0 0.0 0.0 100.0 0.0 $7.3\times 10^{-19}$
 Lentz wall structure is $44\times$ under-resolved at $50^{3}$ (wall-resolution assessment; see Table 3); reported fractions are lower-bound estimates.
 WarpShell uses a $C^{2}$ quintic Hermite regularization of the thin-shell; margins reflect the regularized implementation, not an idealized thin-shell spacetime.
@@ 8 Discussion
Our central finding is that single-frame energy condition evaluation can systematically underestimate the extent of energy condition violations in warp drive spacetimes, though the severity depends strongly on the metric geometry. The Rodal metric has been shown in the original reference [38] to violate the NEC, WEC, DEC, and SEC; our results corroborate those findings under observer-robust analysis and additionally quantify the spatial extent of violations missed by single-frame Eulerian evaluation. For the Rodal metric, the Eulerian DEC analysis fails to detect violations at over 28% of all grid points, and the WEC figure exceeds 15%. For WarpShell (regularized implementation), missed WEC violations are below $0.1\%$ of the grid. For the Alcubierre and Natário metrics, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis at the tested parameters (missed =0=0 at the sampled resolution). Lentz shows zero missed NEC/WEC points (though nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; Table 3); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$, DEC $0.3\%$ at $v_{s}=0.5$; Figures 34), and non-trivial SEC misses persist, most notably $1.2\%$ for Van Den Broeck (conditional miss rate $17.5\%$; Table 4). Even for these metrics the $\zeta_{\max}$-capped violation severity can be much larger (Section 2.4, Interpretation caveat): for the Alcubierre WEC at $\zeta_{\max}=5$, the capped minimum is $\sim\!-3450$ compared to 0.038-0.038 for the Eulerian frame. These severity ratios are $\zeta_{\max}$-dependent and grow without bound; single-frame analysis can underestimate the magnitude of violations even when it correctly identifies their spatial extent.
@@
For the Alcubierre metric at $v_{s}=0.5$, $R_{b}=1$, $\sigma=8$ (Figure 5), we observe that the worst-case WEC observers are predominantly Lorentz-boosted along the direction of bubble propagation. This is physically plausible: boosted observers see a different effective stress-energy distribution, and the Eulerian frame, adapted to the spatial hypersurface, is not aligned with the worst-case direction. For the Rodal DEC, the alignment angle between the optimizer’s boost direction and the DEC-worst eigenvector spatial direction has a median of $82^{\circ}$ across all DEC-violating Type I points at four velocities (Figure 12), with only 0.90.9% of points having alignment below $10^{\circ}$. This confirms that the worst-case DEC observers are overwhelmingly non-Eulerian and that single-frame analysis misses violations precisely because it fails to probe these nearly orthogonal directions. The effect is condition-dependent: the WEC and DEC are more sensitive to observer choice than the NEC, because they involve timelike observers (parameterized by rapidity) rather than null directions alone.
@@
Any claim that a warp metric satisfies an energy condition must specify which observers were tested and with what coverage. The original claim that the Lentz metric admits a positive-energy warp drive was challenged on general theoretical grounds by Santiago, Schuster, and Visser [40], who showed that any physically reasonable warp drive must violate the NEC; this constraint that also implies violation of the WEC, SEC, and DEC. Celmaster and Rubin [8] subsequently demonstrated explicit Eulerian-frame WEC violations in the Lentz geometry, showing that the original positive-energy claim is incorrect even for Eulerian observers. Our observer-robust analysis further confirms SEC violations in the Lentz wall region; NEC/WEC fractions are reported as lower bounds owing to the $50^{3}$ under-resolution of the wall. More generally, our results for Rodal and WarpShell demonstrate that single-frame analysis can miss substantial violation fractions, while for other metrics the Eulerian frame detects all violation points. Continuous observer optimization, as implemented in warpax, quantifies both the spatial extent of missed violations and the severity increase for non-Eulerian observers.
@@
WarpFactory [21, 20, 2] samples $\sim\!1000$ observer directions and velocities at each grid point and takes the minimum energy condition value over the sample set. This discrete-sampling approach correctly identifies observer dependence, but its accuracy depends on the sampling density: narrow violation cones in the observer manifold may be missed if no sample falls within them. A Fibonacci-lattice benchmark (Figure 13) confirms this quantitatively: even $10^{4}$ deterministic samples per point saturate at $\sim\!93$% DEC detection for the Rodal metric, with a sampled minimum margin three orders of magnitude looser than the observer-optimized result. warpax complements sampling with continuous observer optimization (Section 2.4), which targets the worst-case observer via gradient descent with multi-start initialization. The two approaches are complementary: WarpFactory provides a well-tested MATLAB implementation for rapid exploration, while warpax provides gradient-based observer optimization, autodiff curvature, and Hawking–Ellis classification in a GPU-accelerated Python framework. Our Eulerian-baseline NEC/WEC results match WarpFactory’s published values for the Alcubierre metric at matched parameters; residual $\sim\!1$% relative differences are consistent with autodiff vs. finite-difference curvature computation.
@@
Observer optimization is the dominant computational cost. For a single metric at $50^{3}$ resolution with $N_{\text{starts}}=8$ optimizer runs per point, the full analysis takes on the order of minutes on GPU hardware (compared to seconds for the Eulerian-only analysis). For Type I points, the Hawking–Ellis classification provides exact algebraic slacks (for the implemented regularized metric) that bypass the optimization entirely for violation detection; the optimizer cost is then justified primarily by worst-observer identification.
@@
We additionally note that dense sampling at $10^{4}$ random directions detects 100% of the violations that the optimizer finds for the Rodal NEC and WEC. However, for the Rodal DEC, sampling plateaus at only 54% detection even at $10^{4}$ directions, while in our tests the optimizer recovered 100% of the algebraic Type I violation set, a clear case where the violation cone in observer space is too narrow for random sampling to cover reliably. These results demonstrate that while dense sampling suffices for NEC and WEC in most cases, gradient-based optimization substantially outperforms random sampling for DEC analysis, where the violation cone in observer space can be too narrow for random sampling to cover reliably. For Type I points, the algebraic eigenvalue check provides exact DEC truth (for the implemented regularized metric) without any observer search; the optimization advantage is relevant for analyses that rely on observer sampling. Table 11 summarizes these detection rates.
@@
Our analysis is restricted to classical energy conditions. Quantum energy inequalities [14, 34] provide weaker but more physical constraints on the stress-energy, and their systematic evaluation over observers remains an open problem. Additionally, our metrics are parameterized by a finite set of shape parameters ($v_{s}$, $R_{b}$, $\sigma$), and the results may not generalize to metrics outside this family.
@@
Several extensions are natural: (i) metric-space optimization, where the shape function $f(r_{s})$ itself is optimized to minimize energy condition violations; (ii) integration of quantum inequality bounds [14]; (iii) application to the broader family of warp drive metrics, including the Bobrick–Martire classification [6]; and (iv) extension to non-vacuum backgrounds (e.g., cosmological spacetimes).
@@ 9 Conclusion
This work is financially supported by VinUniversity under the Environmental Intelligence (CEI) Grant (No. VUNI.CEI.FS_0009). The author thanks J. Rodal for clarifications regarding frame conventions, the $r_{s}\to 0$ limit of the irrotational angular profile, and the Cartesian form of the shift vector. The author declares no competing interests.
@@ Appendix D Metric definitions
$f(r_{s})=\frac{\tanh\bigl[\sigma(r_{s}+R_{b})\bigr]-\tanh\bigl[\sigma(r_{s}-R_{b})\bigr]}{2\,\tanh(\sigma R_{b})}\,$, (27)
with $f(0)=1$ and $f(r_{s})\to 0$ for $r_{s}\gg R_{b}$.
@@
$G(r_{s})=1-\frac{2\,r_{s}\sigma\sinh(R_{b}\sigma)+\cosh(R_{b}\sigma)\bigl[\ln\cosh\sigma(r_{s}{-}R_{b})-\ln\cosh\sigma(r_{s}{+}R_{b})\bigr]}{2\,r_{s}\sigma\sinh(R_{b}\sigma)}\,.$ (31)
@@
Although equation (31) contains an apparent $1/r_{s}$ factor, this is a removable $0/0$ form arising from the spherical-basis representation. Defining $\Delta(r_{s})\equiv\ln\cosh\bigl[\sigma(r_{s}{-}R_{b})\bigr]-\ln\cosh\bigl[\sigma(r_{s}{+}R_{b})\bigr]$, one finds $\Delta^{\prime}(0)=-2\sigma\tanh(\sigma R_{b})$, whence $\lim_{r_{s}\to 0}G(r_{s})=1$. The irrotational construction is therefore regular at the bubble center; the $r_{s}\mapsto\sqrt{r_{s}^{2}+\varepsilon^{2}}$ substitution used in our implementation is a numerical stability device for automatic differentiation, not a physical repair.